OpenStudy (anonymous):

(2(43.0)+ 2(3.74)^2/1.62) +- sqrt ((2(43.0) + 2(3.74)^2/3.74))^2 - 4(43.0)^2).....i have solved this thing 7 times now and am making errors somewhere...can someone help me?

OpenStudy (anonymous):

it's just the quadratic equation

OpenStudy (anonymous):

oh...the entire equation is divided by 2...smh

OpenStudy (anonymous):

What's the original problem?

OpenStudy (anonymous):

2 trains are on a straight horizontal track. One car starts at rest and is put in motion with a constant acceleration of 1.62 m^2. This car moves toward a car that is 43.0m away and moving at a constant speed of 3.74 m/s. Where will the trains collide?

OpenStudy (anonymous):

So for train 1 we have a function: Distance = 1.62t^2 train 2 is Distance = 43 + 3.74t

OpenStudy (amistre64):

2 trains are on a straight horizontal track. One car starts at rest and is put in motion with a constant acceleration of 1.62 m^2. moves toward a car that is 43.0m away and moving at a constant speed of 3.74 m/s. Where will the trains collide? both cars moving?

OpenStudy (anonymous):

both cars moving

OpenStudy (amistre64):



OpenStudy (anonymous):

phi...in the equation only the numerator is multiplied by 2 not the entire fraction

OpenStudy (anonymous):

i think their moving toward each other amistre.....i'm not sure...question seems vague

OpenStudy (amistre64):

question is very vague in that regards

OpenStudy (anonymous):

If you set the two equations I have above equal to eachother and solve you get t = 6.43406 seconds which is at 67.0633 meters

OpenStudy (amistre64):

a constant acceleration = 1.62 m^2. per sec? per minute?

OpenStudy (anonymous):

i guess b/c it says the trains collide their moving toward each other

OpenStudy (anonymous):

I assumed that they were moving in the same direction

OpenStudy (anonymous):

says 67.1 is incorrect.....we have to use 3 sig figures

OpenStudy (amistre64):

they can be moving in the same direction and the other catch up to it

OpenStudy (amistre64):

a(t) = 1.62 v(t) = 1.62 t + C , but at t=0, v(t)=0, so C = 0 if im thinking right

OpenStudy (anonymous):

If they are moving at eachother change the middle term in the train 2 to a negative. Setting them equal and solving gives time 4.12542 seconds and distance 27.5709 meters

OpenStudy (anonymous):

My equations I'm using are For train 1 we have a function: Distance = 1.62t^2 Train 2 is Distance = 43 - 3.74t

OpenStudy (anonymous):

i took too many tries....answer turned out to be 80.2

OpenStudy (anonymous):

i do not see how that happens

OpenStudy (amistre64):

A train speed = 1.62 t position = 1.62 t^2/ 2 B train speed = 3.74 m/s. position = 3.74 t right?

OpenStudy (phi):

with both trains moving in the same direction http://www.wolframalpha.com/input/?i=solve+0.81+t%5E2+-3.74*t-43%3D0 t= 9.95 seconds distance will be 0.81*t^2

OpenStudy (phi):

using d= 0.5 a t^2 and d= 43 + 3.74 t

OpenStudy (phi):

if the trains move towards each other t= 5.33 seconds d= 23 m

OpenStudy (amistre64):

.81 t^2 = 3.74 t + 40 .81 t^2 -3.74t -40 = 0 t = \(\cfrac{1}{1.62}\left(3.74+\sqrt{3.74^2-4(.81)(-40)}\right)\) = \(\cfrac{1}{1.62}\left(3.74+\sqrt{13.9876+129.6}\right)\) = \(\cfrac{1}{1.62}\left(3.74+\sqrt{143.5876}\right)\) = \(\cfrac{1}{1.62}(15.7228)\) = \(9.7054321\) ............................ B train position = 3.74 t +40 = 3.74(9.7054321) +40 = 76.298... which is pretty close to what they say the answer is

OpenStudy (amistre64):

prolly accounting for my truncating

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i wish i could use the excuse that it's hard...but tis' not....just tedious

OpenStudy (phi):

use 43 instead of 40

OpenStudy (amistre64):

well, and maybe if i used 43 instead of 40 lol

OpenStudy (amistre64):

wolfram says the correction to 43 makes t = 9.95169 and 3.74(9.95169) +43 = 80.219 ...